OpenStudy (anonymous):

i'll give out a metal, help please? Jodie is making bracelets. She measured the length of the bracelets she made for her friends and recorded the measurements in the line plot below. (will share in comments).. She decided to make matching mother and daughter bracelets by doubling the lengths of what she has already made. What will be the length of the longest bracelets for the mothers?

OpenStudy (anonymous):

The largest bracelet in the chart is 3 3/4 inches. Since she is going to double the length, we need to multiply 3 3/4 inches by 2. \[3\frac{ 3 }{ 4 }*2\] Change the mixed number to an improper fraction. Multiple the bottom number times the whole number and then add it to the top number. Keep the bottom number the same. \[3*4=12\]\[3+12=15\]\[\frac{ 15 }{ 4 }\]Now multiply the improper fraction by 2. \[\frac{ 15 }{ 4 }*2=\]Change 2 to a fraction.\[\frac{ 15 }{ 4 }*\frac{ 2 }{ 1 }=\frac{ 15*2 }{ 4*1 }\]\[\frac{ 30 }{ 4 }\]Change improper fraction 30/4 to a mixed number by dividing the bottom number into the top number.\[7\frac{ 2 }{ 4 }\]Reduce the fraction 2/4 \[7\frac{ 1 }{ 2 }inches\]